The base and height of a triangle are (x+3)cm and (2x-5)cm respectively.If the area of the triangle is 20cm squared ,find x
Answer by stanbon(75887)

The base and height of a triangle are (x+3)cm and (2x-5)cm respectively.If the area of the triangle is 20cm squared ,find x
-----
Area = (1/2)(base)(height)
20 = (1/2)(x+3)(2x-5)
40 = 2x^2 +x - 15
2x^2 + x - 55 = 0
------
x = [-1 +- sqrt(1-4*2*(-55))]/4
------
x = [-1 +- sqrt(520)]/4
---
x = 20/4 = 5
----
Cheers,
Stan H.
----------